OpenStudy (anonymous):

Solve for z: x+2y+z= 4 x+2z=3 -y-4z=-5

OpenStudy (anonymous):

you've gotta set up two problems for this one

OpenStudy (anonymous):

@lowcard2 how do i do that?

OpenStudy (***[isuru]***):

Hi, then let's number the equations first, x + 2y + z = 4 ---(1) x + 2z = 3 ---(2) -y - 4z = -5 ---(3) let's multiply (3) by -2 2y + 8z = 10 ---(4) (2) + (4) x + 2y + 10z =13 ---(5) (5) - (1) 9z = 9 z = 1 hope this will help u out

OpenStudy (anonymous):

thank you so much @***[ISURU]*** !